# Out-of-equlibrium inference of feeding rates through population data from generic consumer-resource stochastic dynamics

### v2

@@ 5 Discussion
In this contribution we have studied model identifiability in two stochastic population models, and we have shown that, when independent model observations are taken at different sampling times in the transient regime, the model is truly identifiable in the sense that model parameters can be uniquely estimated from observations. On the contrary, if population numbers are observed at stationarity, then the model is set-identifiable because only a combination of parameters can be inferred from data. We derived model redundancy conditions using the generation function formalism in both a single-popultion birth-death-immigration model, and a multi-resource consumer-resource stochastic population model. For the latter, synthetic data were used to calibrate the model in two in silico experiments: one in which only free predator population numbers are considered, and another in which the whole matrix of searching and handling predator numbers is available. If this model was to be calibrated using actual, experimental data, and independent observations were taken at stationarity (or at equal sampling times), the problem of parameter redundancy would persist: the log-likelihood function would produce an infinite set of maxima, because the critical points conditions would reduce to one single equation —see the derivation of Eq. (99) in B.
@@
Additional information should be available to conduct precise parameter inference at stationarity. When we take logarithms of Eq. (74), and change the sign, we obtain a log-likelihood function, which exactly corresponds to the one represented in the heatmaps of Figure 2 as the two parameters vary between a minimum and a maximum value along the axes. As shown in Section 3, the final expression for the joint stationary distribution (as well as its marginals) depends on the compound parameter controlled by the dimensionless ratio $\nu/\alpha$ (the relaxation rate in units of the attack rate). This non-dimensional quotient is the one that can be statistically inferred from data at equilibrium. If we were able to sort out, experimentally, individuals in the different behavioral modes at stationarity (attacking vs handling), we could identify, as shown in Figure 2, the true attack rate $\alpha$ on resources, respectively, provided we had external information to previously determine the relaxation rate $\nu$, and confidence intervals could be assigned (see Table 1). Regarding to the rest of parameters, this is, the total number of consumers, $n_{A}^{0}$, resource densitities $x_{R_{i}}$, and the volume, $N$, of the system, we believe they will be rather easy to be initially fixed and controlled in the setting of any experiment.
@@
where the second reaction holds for $i\neq j$. This way, depending on the balance between $\alpha_{ii}$ and $\alpha_{ij}$, consumer’s feeding rate on resource $R_{i}$ can be either enhanced or damped by the presence of an alternative resource $R_{j}$ more than it is by the presence of the same type of resource $R_{i}$. Therefore, consumer efficiency on a focal resource is affected not only by the density of the focal resource, but also by the presence of alternative resources.
Proceeding in the same way as in Palamara et al. (2022), the functional response that arises from those fundamental processes corresponds to a multispecies version of Holling type III functional response. It is easy to see that, according to the reaction scheme (80)–(82) and because of mass action, the per capita consumption rate of a free/searching individual $X_{A}$ feeding on a focal resource $i$ becomes
@@
We used the assumption of equal handling times in order to obtain closed-form solutions for the out-of-equilibrium joint distribution of population abundances. By continuity, we expect the multinomial distribution to approximately hold when the distribution of handling times is not broad. In fact, we can always write $\nu_{j}=\nu+\epsilon e_{j}$ (for constant $\nu$), substitute this expression into Eq. (53) and neglect the term $\epsilon\sum_{j=1}^{S}\theta_{j}(1-z_{j})e_{j}$ if we assume that parameter $\epsilon$ is small enough compared to $\nu$. In principle, a perturbation series in this small parameter $\epsilon$ could be carried out in order to solve Eq. (54) for non-uniform rates $\nu_{j}$. Indeed, the closed-form solution we obtained in Section 3 for the PDE is the zero-th order term in the perturbation series expansion on $\epsilon$ of the general solution for non-equal relaxation rates. Future analytical work is needed to provide meaningful expressions for out-of-equilibrium distributions also when the differences in handling times are significant. On a side note, it is important to observe that the assumption of equal handling times does not impose any restriction in the case of one single resource: for the standard Holling type II consumer-resource feeding dynamics, the calculations that follow are fully general for the case $S=1$, and the multinomial out-of-equilibrium solution is exact in that case —provided that the dynamics starts with $n_{A}^{0}$ searching consumers initially.
@@ Acknowledgements
Appendix A Derivation of the PDE satisfied by the generating function
@@
$\frac{\partial G}{\partial t}=A+B,$ (84)
@@
$A:=\sum_{n_{1},\dots,n_{S}\geq 0}\sum_{i=1}^{S}z_{1}^{n_{1}}\cdots z_{S}^{n_{S}}\left[T(\bm{n}|\bm{n}+\bm{e_{i}})P_{\bm{n}+\bm{e_{i}}}-T(\bm{n}-\bm{e_{i}}|\bm{n})P_{\bm{n}}\right]$ (85)
@@
$B:=\sum_{n_{1},\dots,n_{S}\geq 0}\sum_{i=1}^{S}z_{1}^{n_{1}}\cdots z_{S}^{n_{S}}\left[T(\bm{n}|\bm{n}-\bm{e_{i}})P_{\bm{n}-\bm{e_{i}}}-T(\bm{n}+\bm{e_{i}}|\bm{n})P_{\bm{n}}\right],$ (86)
@@
$\frac{\partial G}{\partial z_{i}}=\sum_{n_{1},\dots,n_{S}\geq 0}(n_{i}+1)P_{\bm{n}+\bm{e}_{i}}z_{1}^{n_{1}}\cdots z_{S}^{n_{S}},$ (87)
@@
$A=\sum_{n_{1},\dots,n_{S}\geq 0}\sum_{i=1}^{S}[\nu_{i}(n_{i}+1)P_{\bm{n}+\bm{e}_{i}}-\nu_{i}n_{i}P_{\bm{n}}]z_{1}^{n_{1}}\cdots z_{S}^{n_{S}}=\sum_{i=1}^{S}\nu_{i}(1-z_{i})\frac{\partial G}{\partial z_{i}}.$ (89)
@@
$B=\sum_{n_{1},\dots,n_{S}\geq 0}\sum_{i=1}^{S}z_{1}^{n_{1}}\cdots z_{S}^{n_{S}}\biggl{\{}n_{A}^{0}\theta_{i}\left(P_{\bm{n}-\bm{e_{i}}}-P_{\bm{n}}\right)+\theta_{i}\biggl{[}P_{\bm{n}}\sum_{j}n_{j}-\biggl{(}n_{i}-1+\sum_{j\neq i}n_{j}\biggr{)}P_{\bm{n}-\bm{e}_{i}}\biggr{]}\biggr{\}}.$ (90)
@@
$z_{i}^{2}\frac{\partial G}{\partial z_{i}}$ $=\sum_{n_{1},\dots,n_{S}\geq 0}(n_{i}-1)P_{\bm{n}-\bm{e}_{i}}z_{1}^{n_{1}}\cdots z_{S}^{n_{S}},$ (91)
@@
$B=\sum_{i=1}^{S}\theta_{i}(1-z_{i})\biggl{(}-n_{A}^{0}G+\sum_{j=1}^{S}z_{j}\frac{\partial G}{\partial z_{j}}\biggr{)}.$ (93)
@@
Appendix B ML estimation with equal sampling times
@@
$P_{\bm{n}}(n_{A}^{0},\{p_{i}\})=C(\bm{n})\left(1-\sum_{i=1}^{S}p_{i}\right)^{n_{A}}\prod_{i=1}^{S}p_{i}^{n_{i}},$ (94)
@@
$p_{j}(t)=\frac{x_{R_{j}}\alpha}{\nu}\cdot\frac{1-e^{-\left(\nu+\alpha\Lambda_{R}\right)t}}{1+\alpha\Lambda_{R}/\nu}.$ (95)
@@
$\log\mathcal{L}(\alpha,\nu|M(t))=T\log C(\bm{n})+\sum_{k=1}^{T}\left\{n_{AR,k}\log h(t;\alpha,\tau)+n_{A,k}\log(1-\Lambda_{R}h(t;\alpha,\tau))+\sum_{i=1}^{S}n_{i,k}\log x_{R_{i}}\right\},$ (97)
@@
$\frac{\partial}{\partial\alpha}\log\mathcal{L}=\frac{\partial h}{\partial\alpha}\sum_{k=1}^{T}\left(\frac{n_{AR,k}}{h}-\frac{n_{A,k}\Lambda_{R}}{1-\Lambda_{R}h}\right)$ $=0,$ (98)
@@
$\sum_{k=1}^{T}\left(\frac{n_{AR,k}}{h}-\frac{n_{A,k}\Lambda_{R}}{1-\Lambda_{R}h}\right)=0,$ (99)
@@
where we have defined $\langle n_{AR}\rangle_{T}:=\frac{1}{T}\sum_{k=1}^{T}n_{AR,k}$ as the average across samples of handling consumers’ total abundance. The quantity $\langle n_{AR}\rangle_{T}$, relative to the product $\Lambda_{R}n_{A}^{0}$, determines only one single restriction to be satisfied by model parameters, in the $(\alpha,\nu)$ space. This immediately implies that the two parameters cannot be uniquely determined, yielding an infinite number of critical points satisfying Eq. (100), as we wanted to prove. In the case of exactly equal sampling times (either out of equilibrium or at stationarity), the model remains set-identifiable. At least two sampling times must be distinct in order for $\alpha$ and $\nu$ being non-redundant.
@@
$\frac{\nu}{\alpha}=\frac{\langle n_{A}\rangle_{T}}{\langle n_{AR}\rangle_{T}}\Lambda_{R},$ (101)

### v3

@@ 5 Discussion
In this contribution we have studied model identifiability in two stochastic population models, and we have shown that, when independent model observations include samples taken at different sampling times in the transient regime, the model is truly identifiable in the sense that model parameters can be uniquely estimated from observations. On the contrary, if population numbers are only observed at stationarity, then the model is set-identifiable because only a combination of parameters can be inferred from data. We derived model redundancy conditions using the generation function formalism in both a single-popultion birth-death-immigration model, and a multi-resource consumer-resource stochastic population model. For the latter, synthetic data were used to calibrate the model in two in silico experiments: one in which only free predator population numbers are considered, and another in which the whole matrix of searching and handling predator numbers is available. If this model was to be calibrated using actual, experimental data, and independent observations were taken at stationarity (or at equal sampling times), the problem of parameter redundancy would persist: the log-likelihood function would produce an infinite set of maxima, because the critical points conditions would reduce to one single equation —see the derivation of Eq. (49) in C.
Importantly, if temporal samples contain both transient and stationary observations, model parameters remain perfectly identifiable. Therefore, an important implication of our work is that time series must include out-of-equilibrium observations to infer parameters without redundancy. Several statistical methods allow one to separate transient and steady-state regimes in a time series: for example, the Chow test (Chow, 1960) allows to separate a positive slope in abundance time series from a slope compatible with zero, proper of a steady-state regime. For empirical data, the experiment has to be designed to collect out-of-equilibrium samples, by monitoring the experiment from an initial condition far from equilibrium.
@@
Additional information should be available to conduct precise parameter inference at stationarity. The heatmaps in Fig. 2 represent the log-likelihood function values, Eq. (30), as the two parameters vary between a minimum and a maximum value along the axes, and show that when samples are taken close enough to stationarity, then a line of optimal values arise. This agrees with the expression found in A for the joint stationary distribution (as well as its marginals), which depends on the compound parameter controlled by the dimensionless ratio $\nu/\alpha$ (the relaxation rate in units of the attack rate). This nondimensional quotient is the one that can be statistically inferred from data at equilibrium. If we were able to sort out, experimentally, individuals in the different behavioral modes at stationarity (attacking vs handling), we could identify, as shown in Figure 2, the true attack rate $\alpha$ on resources, respectively, provided that we had external information to previously determine the relaxation rate $\nu$, and confidence intervals could be assigned (see Table 1). Regarding to the rest of parameters, that is, the total number of consumers, $n_{A}^{0}$, resource densities $x_{R_{i}}$, and the volume, $N$, of the system, we believe that they will be fairly easy to initially fix and control in the setting of any experiment.
We stress that any additional information used to solve parameter redundancy at equilibrium should come from monitoring the system out of equilibrium. In particular, precise estimates of $\tau$ would break redundancy at stationarity. Observe that, if the characteristic time $\tau$ and the equilibrium distribution are known, one can obtain non-redundant parameter estimates in the case of equal handling times. However, estimating the characteristic time requires knowledge of the transient dynamics before reaching the steady state. One could perturb an experimental system by varying populations along a proper direction away from the equilibrium point (along an eigenvector of the Jacobian matrix evaluated at equilibrium). Then the eigenvalue would help measure the relaxation time. Methods alternative to our ML approach, such as the one proposed in D (based on the initial decay of the abundance of free predators), imply breaking stationarity or having precise knowledge of the transient state. These methods are, therefore, equivalent to sampling the system during the transient dynamics before stationarity.
@@
where the second reaction holds for $i\neq j$. In this way, depending on the balance between $\alpha_{ii}$ and $\alpha_{ij}$, consumer’s feeding rate on resource $R_{i}$ can be enhanced or damped by the presence of an alternative resource $R_{j}$ more than it is by the presence of the same type of resource $R_{i}$. Therefore, consumer efficiency on a focal resource is affected not only by the density of the focal resource but also by the presence of alternative resources.
Proceeding in the same way as in Palamara et al. (2022), the functional response that arises from those fundamental processes corresponds to a multispecies version of the Holling type III functional response. It is easy to see that, according to the reaction scheme (33)–(35) and because of mass action, the per capita consumption rate of a free/searching individual $X_{A}$ feeding on a focal resource $i$ becomes
@@
We used the assumption of equal handling times to obtain closed-form solutions for the out-of-equilibrium joint distribution of population abundances. When handling times for different resources are not equal, our methodology applies if the master equation is integrated numerically to obtain probability distributions as functions of time. Given these probabilities, one can compute numerically log-likelihoods and estimate the $2S$ parameters $\{\alpha_{i},\nu_{i}\}$ ($i=1,\dots,S$) without redundancy. It is important to note that, in this case, the degeneracy of the approach when samples are taken at equilibrium increases because, in this case, only $S$ ratios ($\alpha_{i}/\nu_{i}$) are identifiable —see Eqs. (12)–(13). If only abundance data at stationarity is available, to solve redundancy one would need prior knowledge of $S$ additional quantities relating the unknown rates. On a side note, it is important to observe that the assumption of equal handling times does not impose any restriction in the case of one single resource: for the standard Holling type II consumer-resource feeding dynamics, the calculations that follow are fully general for the case $S=1$, and the multinomial out-of-equilibrium solution is exact in that case —provided that the dynamics starts with $n_{A}^{0}$ searching consumers initially.
Traditional methods for estimating attack and handling rates in consumer-resource interactions rely on depletion experiments, where a fixed number of consumers are placed with a known initial resource density and the abundance of resources decreases over time due to consumption (Rogers, 1972). Parameters like attack rate and handling time are inferred by fitting the depletion curve to a model of resource abundance. These experiments run out-of-equilibrium, without resource replenishment. In contrast, our approach assumes chemostatic conditions, where consumed resources are continuously replenished, and the total number of consumers remains constant. By tracking consumer behavior —distinguishing between searching and handling states— through animal trajectories, one could apply our maximum likelihood approach to estimate process rates. Monitoring the early stages of the experiment is crucial to avoid parameter redundancy. Experimental setups capable of classifying individual behaviors, such as searching for or digesting resources, can support this approach and allow parameter estimation under controlled conditions (Bartumeus et al., 2002, 2005, 2003). Because maximum likelihood estimators are asymptotically those with minimal variance, we expect that our approach is better than simply fitting a curve to data, as in depletion experiments.
@@ Acknowledgements
Code availability
Code to reproduce all findings is available at https://github.com/vankampen92/PROJECT_HOLLING_II_nD
Appendix A Steady-state distribution derivation
$P_{n_{A}}=\binom{n_{A}^{0}}{n_{A}}p^{n_{A}}(1-p)^{n_{A}^{0}-n_{A}},$ (37)
$n_{A}$ being the number of searching consumers (Palamara et al., 2022). One can take advantage of knowing that distribution to rigorously derive the functional response associated to the single resource case ($S=1$), which turns out to be the Holling type II functional response (Palamara et al., 2022). We here extend the binomial distribution result to a multi-resource case by solving the master equation at stationarity, which yields an analytical expression for the joint steady-state distribution $P_{\bm{n}}$. We can do this because transition probability rates (9) are linear in population numbers (Kendall, 1949; Nisbet and Gurney, 2004). As a result, the single-resource binomial distribution is extended to a multi-resource case, leading to a multinomial distribution for the joint probability distribution of observing a vector $(n_{A},\bm{n})$ of the populations of behavioral types, as we show next.
$T(\bm{n}-\bm{e_{i}}|\bm{n})P_{\bm{n}}=T(\bm{n}|\bm{n}-\bm{e_{i}})P_{\bm{n}-\bm{e_{i}}},$
and, subsequently,
$P_{\bm{n}+\bm{e}_{i}}=\frac{\nu_{j}}{\theta_{j}}\frac{\theta_{i}}{\nu_{i}}\frac{n_{j}+1}{n_{i}+1}P_{\bm{n}+\bm{e}_{j}}.$
Making the change $n_{i}\to n_{i}-1$ in the last equation one gets
Now, apply Eq. (41) to resources $i=1$ and $j=2$: this gives
$P_{\bm{n}}=\frac{\theta_{1}}{\nu_{1}}\frac{\nu_{2}}{\theta_{2}}\frac{n_{2}+1}{n_{1}}P_{(n_{1}-1,n_{2}+1,\dots,n_{S})}.$
$P_{\bm{n}}=\left(\frac{\theta_{1}}{\nu_{1}}\frac{\nu_{2}}{\theta_{2}}\right)^{2}\frac{(n_{2}+1)(n_{2}+2)}{n_{1}(n_{1}-1)}P_{(n_{1}-2,n_{2}+2,\dots,n_{S})}.$
By repeating the iteration $n_{1}-2$ additional times one gets
$P_{\bm{n}}=\left(\frac{\theta_{1}}{\nu_{1}}\frac{\nu_{2}}{\theta_{2}}\right)^{n_{1}}\binom{n_{1}+n_{2}}{n_{1}}P_{(0,n_{1}+n_{2},\dots,n_{S})},$
$P_{\bm{n}}=\left(\frac{\theta_{1}}{\nu_{1}}\frac{\nu_{2}}{\theta_{2}}\right)^{n_{1}}\left(\frac{\theta_{2}}{\nu_{2}}\frac{\nu_{3}}{\theta_{3}}\right)^{n_{1}+n_{2}}\binom{n_{1}+n_{2}}{n_{1}}\binom{n_{1}+n_{2}+n_{3}}{n_{1}+n_{2}}P_{(0,0,n_{1}+n_{2}+n_{3},\dots,n_{S})},$
which simplifies to
$P_{\bm{n}}=\left(\frac{\theta_{1}}{\nu_{1}}\right)^{n_{1}}\left(\frac{\theta_{2}}{\nu_{2}}\right)^{n_{2}}\left(\frac{\nu_{3}}{\theta_{3}}\right)^{n_{1}+n_{2}}\binom{n_{1}+n_{2}+n_{3}}{n_{1},n_{2},n_{3}}P_{(0,0,n_{1}+n_{2}+n_{3},\dots,n_{S})},$
$P_{\bm{n}-\bm{e}_{S}}=\frac{\nu_{S}n_{S}}{\theta_{1}(n_{A}^{0}+1-\sum_{k=1}^{S}n_{k})}P_{\bm{n}}.$
$P_{(q-1)\bm{e}_{S}}$ $=\frac{\nu_{S}}{\theta_{S}}\frac{q}{n_{A}^{0}+1-q}P_{q\bm{e}_{S}},$
[…]
$P_{q\bm{e}_{S}}=\binom{n_{A}^{0}}{q}\left(\frac{\theta_{S}}{\nu_{S}}\right)^{q}P_{\bm{0}}.$
where we have used Eq. (8) and
$\binom{n_{A}^{0}}{\sum_{k=1}^{S}n_{k}}\binom{\sum_{k=1}^{S}n_{k}}{n_{1},\dots,n_{S}}=\frac{n_{A}^{0}!}{n_{1}!\cdots n_{S}!n_{A}!}.$
In the form of Eq. (43), the joint stationary distribution remarkably resembles a multinomial distribution, see Eq. (11). To write Eq. (43) as a multinomial distribution, first we have to calculate the normalization constant $P_{\bm{0}}$. It follows immediately after evaluating the probability of observing a total number $q$ of handling predators, $q=\sum_{k=1}^{S}n_{k}$, which can be done as follows.
$P_{q}=\sum_{\sum_{i}n_{i}=q}\frac{n_{A}^{0}!}{n_{1}!\cdots n_{S}!n_{A}!}\left[\prod_{k=1}^{S}\left(\frac{\theta_{k}}{\nu_{k}}\right)^{n_{k}}\right]P_{\bm{0}},$
which can be written alternatively as
$P_{q}=\binom{n_{A}^{0}}{q}\sum_{\sum_{i}n_{i}=q}\frac{q!}{n_{1}!\cdots n_{S}!}\left[\prod_{k=1}^{S}\left(\frac{\theta_{k}}{\nu_{k}}\right)^{n_{k}}\right]P_{\bm{0}}.$
$\sum_{\sum_{i}n_{i}=q}\frac{q!}{n_{1}!\cdots n_{S}!}\prod_{k=1}^{S}\left(\frac{\theta_{k}}{\nu_{k}}\right)^{n_{k}}=\left(\sum_{k=1}^{S}\frac{\theta_{k}}{\nu_{k}}\right)^{q},$
which implies that
Therefore, if we impose normalization of the distribution $P_{q}$ we get
where in the last equality we have used the binomial theorem.
The joint distribution of the vector of population abundances of different behavioral types, $(n_{A},n_{1},\dots,n_{S})$, is given by Eqs. (43) and (45). It is a multinomial distribution $\mathcal{M}(n_{A}^{0},\{p_{A},p_{1},\dots,p_{S}\})$, as defined in Eq. (11), with $n_{A}^{0}$ trials and probabilities given by Eqs. (12) and (13), which we reproduce here:
$p_{i}=\frac{\theta_{i}}{\nu_{i}}\frac{1}{1+\sum_{k=1}^{S}\frac{\theta_{k}}{\nu_{k}}}$
for predators handling on resources $i=1,\dots,S$, and
$p_{A}=\frac{1}{1+\sum_{k=1}^{S}\frac{\theta_{k}}{\nu_{k}}}$
for free predators. Observe that $p_{A}+\sum_{k=1}^{S}p_{k}=1$ is satisfied. In addition, Eq. (13) extends naturally, for more than one resource, the probability parameter $p=\left(1+\frac{\theta}{\nu}\right)^{-1}$ of the binomial distribution (37).
and a mean number of free predators equal to
In addition to the joint, steady-state probability distribution of different behavioral types, by summing over the corresponding variables we can obtain the marginal distributions of each single type, which are binomials $B(n_{A}^{0},p_{i})$ (Loève, 1977). In particular, from our derivation above we find that the distribution of the total number of handling consumers is a binomial $B(n_{A}^{0},1-p_{A})$: using Eqs. (45) and (13), the probability that $\sum_{k=1}^{S}n_{k}=q$, which is given by Eq. (44), can be rewritten as
$P_{q}=\binom{n_{A}^{0}}{q}(1-p_{A})^{q}p_{A}^{n_{A}^{0}-q}.$
$P_{n_{A}}=\binom{n_{A}^{0}}{n_{A}}p_{A}^{n_{A}}(1-p_{A})^{n_{A}^{0}-n_{A}},$
as expected (Palamara et al., 2022).
Appendix B Derivation of the PDE satisfied by the generating function
@@
$\frac{\partial G}{\partial t}=A+B,$
@@
$A:=\sum_{n_{1},\dots,n_{S}\geq 0}\sum_{i=1}^{S}z_{1}^{n_{1}}\cdots z_{S}^{n_{S}}\left[T(\bm{n}|\bm{n}+\bm{e_{i}})P_{\bm{n}+\bm{e_{i}}}-T(\bm{n}-\bm{e_{i}}|\bm{n})P_{\bm{n}}\right]$
@@
$B:=\sum_{n_{1},\dots,n_{S}\geq 0}\sum_{i=1}^{S}z_{1}^{n_{1}}\cdots z_{S}^{n_{S}}\left[T(\bm{n}|\bm{n}-\bm{e_{i}})P_{\bm{n}-\bm{e_{i}}}-T(\bm{n}+\bm{e_{i}}|\bm{n})P_{\bm{n}}\right],$
@@
$\frac{\partial G}{\partial z_{i}}=\sum_{n_{1},\dots,n_{S}\geq 0}(n_{i}+1)P_{\bm{n}+\bm{e}_{i}}z_{1}^{n_{1}}\cdots z_{S}^{n_{S}},$
@@
$A=\sum_{n_{1},\dots,n_{S}\geq 0}\sum_{i=1}^{S}[\nu_{i}(n_{i}+1)P_{\bm{n}+\bm{e}_{i}}-\nu_{i}n_{i}P_{\bm{n}}]z_{1}^{n_{1}}\cdots z_{S}^{n_{S}}=\sum_{i=1}^{S}\nu_{i}(1-z_{i})\frac{\partial G}{\partial z_{i}}.$
@@
$B=\sum_{n_{1},\dots,n_{S}\geq 0}\sum_{i=1}^{S}z_{1}^{n_{1}}\cdots z_{S}^{n_{S}}\biggl{\{}n_{A}^{0}\theta_{i}\left(P_{\bm{n}-\bm{e_{i}}}-P_{\bm{n}}\right)+\theta_{i}\biggl{[}P_{\bm{n}}\sum_{j}n_{j}-\biggl{(}n_{i}-1+\sum_{j\neq i}n_{j}\biggr{)}P_{\bm{n}-\bm{e}_{i}}\biggr{]}\biggr{\}}.$
@@
$z_{i}^{2}\frac{\partial G}{\partial z_{i}}$ $=\sum_{n_{1},\dots,n_{S}\geq 0}(n_{i}-1)P_{\bm{n}-\bm{e}_{i}}z_{1}^{n_{1}}\cdots z_{S}^{n_{S}},$
@@
$B=\sum_{i=1}^{S}\theta_{i}(1-z_{i})\biggl{(}-n_{A}^{0}G+\sum_{j=1}^{S}z_{j}\frac{\partial G}{\partial z_{j}}\biggr{)}.$
@@
Appendix C ML estimation with equal sampling times
@@
$P_{\bm{n}}(n_{A}^{0},\{p_{i}\})=C(\bm{n})\left(1-\sum_{i=1}^{S}p_{i}\right)^{n_{A}}\prod_{i=1}^{S}p_{i}^{n_{i}},$
@@
$p_{j}(t)=\frac{x_{R_{j}}\alpha}{\nu}\cdot\frac{1-e^{-\left(\nu+\alpha\Lambda_{R}\right)t}}{1+\alpha\Lambda_{R}/\nu}.$
@@
$\log\mathcal{L}(\alpha,\nu|M(t))=T\log C(\bm{n})+\sum_{k=1}^{T}\left\{n_{AR,k}\log h(t;\alpha,\tau)+n_{A,k}\log(1-\Lambda_{R}h(t;\alpha,\tau))+\sum_{i=1}^{S}n_{i,k}\log x_{R_{i}}\right\},$
@@
$\frac{\partial}{\partial\alpha}\log\mathcal{L}=\frac{\partial h}{\partial\alpha}\sum_{k=1}^{T}\left(\frac{n_{AR,k}}{h}-\frac{n_{A,k}\Lambda_{R}}{1-\Lambda_{R}h}\right)$ $=0,$
@@
$\sum_{k=1}^{T}\left(\frac{n_{AR,k}}{h}-\frac{n_{A,k}\Lambda_{R}}{1-\Lambda_{R}h}\right)=0,$
@@
where we have defined $\langle n_{AR}\rangle_{T}:=\frac{1}{T}\sum_{k=1}^{T}n_{AR,k}$ as the average across samples of handling consumers’ total abundance. The quantity $\langle n_{AR}\rangle_{T}$, relative to the product $\Lambda_{R}n_{A}^{0}$, determines only one single restriction to be satisfied by model parameters, in the $(\alpha,\nu)$ space. This immediately implies that the two parameters cannot be uniquely determined, yielding (for every $t>0$) an infinite number of critical points satisfying Eq. (49), as we wanted to prove. In the case of exactly equal sampling times (either out of equilibrium or at stationarity), the model remains set-identifiable. At least two sampling times must be distinct in order for $\alpha$ and $\nu$ being non-redundant.
@@
$\frac{\nu}{\alpha}=\frac{\langle n_{A}\rangle_{T}}{\langle n_{AR}\rangle_{T}}\Lambda_{R},$